An automobile radiator contains 16 quarts of a 20% solution of antifreeze. How much of the original solution must be drawn off and replaced by pure antifreeze to make a 25% antifreeze?
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

An automobile radiator contains 16 quarts of a 20% solution of antifreeze. How much of the original solution must be drawn off and replaced by pure antifreeze to make a 25% antifreeze?
----
Equation:
alcohol - alcohol + alcohol = alcohol
0.20*16 - 0.20x + 1.00x = 0.25*16
----
Multiply thru by 100 to get:
20*16 - 20x + 100x = 25*16
80x = 5*16
x = 5(16/80)
x = 5(1/5)
x = 1 quart (amt. of antifreeze that must be removed and replaced)
==========================
Cheers,
Stan H.

Answer by mananth(16946)

An automobile radiator contains 16 quarts of a 20% solution of antifreeze. How much of the original solution must be drawn off and replaced by pure antifreeze to make a 25% antifreeze?
..
16 quarts of a 20% solution
pure antifreeze 100% x quarts
25% antifreeze mix x+16 quarts
...
concentration equation
16*20+100x=25(x+16)
320+100x=25x+400
100x-25x=400-320
75x=80
/75
x= 80/75
x= 1.06 quarts of antifeeze
...
m.ananth@hotmail.ca